Find two numbers whose sum is 72 and whose difference is 10. The two numbers are?
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52781)

.

Let x be the smaller number.

Then the other number is x+10.


We have this equation


    x + (x+10) = 72


Then


    2x + 10 = 72

    2x = 72 - 10

    2x = 62

     x = 62/2 = 31.


ANSWER.  The numbers are  31  (the smaller)  and  31+10 = 41 (the greater).


Answer by MathTherapy(10552)


Find two numbers whose sum is 72 and whose difference is 10. The two numbers are?
Let larger number be L, and smaller, S
Then we get: L + S = 72 ------- eq (i)
Also, L - S = 10 ------- eq (ii)
Add the 2 equations and solve to get the value of L, the larger number.
Substitute the value of L in any of the 2 equations, preferably eq (i), to get the value of S, the smaller number.